OpenStudy (anonymous):

Calculating and shorten as far as possible. a) 5/9 - 2/3 + 11/18 b) 7/12 / 2/9

OpenStudy (anonymous):

For a) get everything over a common denominatorthat would be 18 multiply 5/9 by 2/2 to get 10/18 multiply -2/3 by 6/6 to get -12/18 add the numerators and leave the denominator alone (10-12+11)/18=9/18=1/2 b) is a complex fraction, so you'll want to multiply by the least common denominator, which would be (9)(12).  7*9=63 and 2*12=24, so now you have 63/24